Fran works due north of home.Her husband Ben works due east. They leave for work at the same time. By the time fran is 3 miles from home, the distance between them is one mile more than bens distance. What is bens distance?
Answer by henew(13)


Draw a picture. You will see a right angled triangle with sides 3,a( ben's distance) and hypotenuse a+1.
By the property of right angled triangle
3 squared + a squared = (a+1) squared.
a^2+2a+1= 9+a^2
2a+1=9
a=4 miles